OpenStudy (anonymous):

What is the inverse function of y=3x-8.7

Vocaloid (vocaloid):

"switch" y and x and solve for y so, if the original function is y = 3x - 8.7, you switch y and x to get x = 3y - 8.7 solve again for y

OpenStudy (anonymous):

To find the inverse, switch the x and y. x = 3y-8 Solve for y. (x+8)/3 = y

OpenStudy (anonymous):

thats the inverse

OpenStudy (anonymous):

Can u explain how u solve for y please

Vocaloid (vocaloid):

sure, so we have: x = 3y - 8.7 what do you get when you add 8.7 to both sides?

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

after you add 8.7 to both sides: x + 8.7 = 3y - 8.7 + 8.7 = 3y, are you with me so far?

Vocaloid (vocaloid):

all I did was add "+ 8.7" to both sides

OpenStudy (anonymous):

Okay

Vocaloid (vocaloid):

now, divide both sides by 3

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So its (x+8.7)/3=y

Vocaloid (vocaloid):

yup! great job!

OpenStudy (anonymous):

Cool thanks